Solve by the substitution method.
  x - 7y = 28
-6x - 6y = 24

Answer by jack.wang5(5)

x - 28 = 7y , so y = x/7 - 4.
- 6x - 6y = 24 , so -6( x + y )=24, x + y = -4 = x + (x/7-4)=-4, x = 0.
y = -4.